# Vim-F: Visual State Space Model Benefiting from Learning in the Frequency Domain

### v1

@@ 4.3 Efficiency analysis
PVTv2-B0 Wang et al., (2022) 3 0.6 70.5
DeiT-Ti Touvron et al., (2021) 6 1.3 74.4
MobileViT-XS Mehta and Rastegari, (2021) 2 1.0 74.8
LVT Yang et al., (2022) 6 0.9 74.8
@@
Vim-Ti-F(H) (ours) 7 1.5 76.0
LocalVim-T Huang et al., (2024) 8 1.5 76.5
RegNetY-1.6G Radosavovic et al., (2020) 11 1.6 78.0
MobileViT-S Mehta and Rastegari, (2021) 6 2.0 78.4
EfficientVMamba-S Pei et al., (2024) 11 1.3 78.7
DeiT-S Touvron et al., (2021) 22 4.6 79.8
RegNetY-4G Radosavovic et al., (2020) 21 4.0 80.0
Swin-T Liu et al., (2021) 29 4.5 81.3
[…]
Vim-S-F(H) (ours) 26 5.1 80.4
LocalVim-SHuang et al., (2024) 28 4.8 81.0
EfficientVMamba-BPei et al., (2024) 33 4.0 81.8
VMamba-T Liu et al., (2024) 22 5.6 82.2
5 Experiments
This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection and instance segmentation.
5.1 Image classification
Settings. We train the models on ImageNet-1K and evaluate the performance on ImageNet-1K validation set. For fair comparisons, our training settings mainly follow Vim. Specifically, we apply random cropping, random horizontal flipping, label-smoothing regularization, mixup, and random erasing as data augmentations. When training on $224\times 224$ input images, we employ AdamW with a momentum of 0.9 and a weight decay of 0.05 to optimize models. During testing, we apply a center crop on the validation set to crop out $224\times 224$ images. We train the Vim-F(H) models for 300 epochs using a cosine schedule and EMA. Unlike Vim, our experiments are performed on 3 A6000 GPUs. Therefore, we adjusted the total batch size and the initial learning rate to 384 and $3.75\times 10^{-4}$ respectively according to the linear scaling ruleGoyal et al., (2017).
Results. We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table 2. Compared to our seminal Vim, our method has made substantial progress, with Vim-Ti-F(H) and Vim-S-F(H) achieving 1.3% and 0.8% higher accuracy than Vim-Ti and Vim-S, respectively. It is worth noting that SOTA ViMs such as VMamba Liu et al., (2024), LocalMamba Huang et al., (2024), and EfficientVMamba Pei et al., (2024) all adopt more complex hybrid encoding architectures, even introducing attention mechanisms. However, Vim-F still uses a pure SSM encoder to model visual information. Vim-Ti-F(H) achieves superior accuracy compared to highly optimized lightweight ViT variants, including DeiT-TiTouvron et al., (2021), MobileViT-XS Mehta and Rastegari, (2021), and LVT Yang et al., (2022), while also narrowing the gap in performance between Vim and other leading ViMs.
Long sequence fine-tuning. ViM with a pure Mamba encoder can compress signals at a low cost and reconstruct them into longer sequences. For visual tasks, flattening images into longer 1D sequences and fine-tuning the model can allow the model to capture more details and improve performance. Due to the use of absolute position embedding, Vim needs to expand the encoding length through interpolation methods. In addition, because the position embedding is located at the front of the entire network, more epochs of fine-tuning are required for adaptation. We set the stride of the first convolutional layer in the patch embedding to 2, and set the learning rate to $1\times 10^{-5}$ for 1 epoch of fine-tuning. Vim-Ti-F(H) can achieve a 0.8% increase in accuracy.
Table 3: The classification result of Vim-F(H) on ImageNet-1K after 1 epoch of fine-tuning. Different from the pre-training stage, the input image is flattened into 784 tokens (28$\times$28).
Method Params (M) Top-1 (%)
Vim-Ti-F(H) (Fine-tuning for 1 epoch) 7 76.8
Vim-S-F(H) (Fine-tuning for 1 epoch) 26 81.2
Table 4: Object detection and instance segmentation results on COCO.
Mask R-CNN 1$\times$ schedule
[…]
LightViT-T 28M 187G 37.8 60.7 40.4 35.9 57.8 38.0
@@ 5.1 Image classification
ConvNeXt-S 70M 348G 45.4 67.9 50.0 41.8 65.2 45.1
Vim-Ti 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
Vim-Ti-F(H) (ours) 27M 189G 37.7 60.8 40.9 35.7 57.8 37.9
Vim-S 44M 272G 42.9 65.4 47.1 39.2 62.1 42.1
Vim-S-F(H) (ours) 44M 272G 43.1 65.2 47.3 39.3 62.2 42.3
5.2 Object detection and instance segmentation
@@
Settings. We use Mask-RCNN as the detector to evaluate the performance of the proposed Vim-F(H) for object detection and instance segmentation on the MSCOCO 2017 dataset. Since the backbone of Vim-F(H) is non-hierarchical, following ViTDetLi et al., (2021), we only used the last feature map from the backbone and generated multi-scale feature maps through a set of convolutions or deconvolutions to adapt to the detector. The remaining settings were consistent with SwinLiu et al., (2021). Specifically, we employ the AdamW optimizer and fine-tune the pre-trained classification models (on ImageNet-1K) for both 12 epochs (1$\times$ schedule). The learning rate is initialized at $1\times 10^{-4}$ and is reduced by a factor of 10×10$\times$10 × at the 9th and 11th epochs.
@@
Results. We summarize the comparison results of Vim-F(H) with other backbones in Table 4. It can be seen that our Vim-F(H) consistently outperforms Vim. Similar to the results on classification tasks, Vim-F(H) with a pure SSM encoder achieves a good balance between the number of parameters and computational cost, achieving comparable results with advanced CNNs and ViTs.
6 Conclusion and limitation
@@
One of the main challenges currently faced by ViMs is the destruction of local two-dimensional dependency relationships when spatial data is flattened into one-dimensional tokens due to the non-causal nature of visual data. Consequently, scanning one-dimensional tokens with specific rules can limit the model’s ability to accurately interpret spatial relationships. We observed that scanning in the frequency domain, leveraging the translation invariance and global nature of Fourier transforms, can reduce the inductive bias of scanning strategies while ensuring that the model always maintains a good global receptive field. We adopted a simple approach of adding the spectrogram with the original feature map, allowing the model to scan simultaneously in both the frequency and spatial domains, thereby establishing a unified representation of visual information across both domains. Furthermore, we employ two trainable parameters to dynamically adjust the information intensity in the frequency and spatial domains, aiming to reduce semantic gaps. We also introduce additional local correlations during patch embedding by using overlapping convolutions. Our Vim-F(H) is motivated by a simple idea and is easy to implement. Extensive experiments have demonstrated that it effectively improves the performance of Vim, achieving competitive results compared to advanced ViTs, CNNs, and ViMs. The effectiveness of our method for ViMs with hybrid encoders has not been fully studied.
Appendix A Appendix
A.1 Ablation study
Effect of frequency domain scanning. We evaluated the scanning strategy of Vim-F, and the experimental details are shown in Table 5. As mentioned earlier, position embeddings are crucial for Vim, and removing them leads to a 1.7% performance drop for Vim-Ti. However, the global receptive field brought by frequency domain scanning reduces the model’s dependence on positional embeddings, resulting in a smaller performance drop for Vim-Ti-F. When both models use positional embeddings, Vim-Ti-F outperforms Vim-Ti by 0.4%.
Effect of convolutional stem. As shown in Table 5, further replacing the traditional patch embedding in Vim-Ti with our proposed convolutional stem leads to a 0.9% improvement in performance. This shows that introducing additional local correlation into the patch embedding is effective.
Table 5: Ablation study of frequency domain scanning and patch embedding in Vim-Ti-F(H) on ImageNet-1K.
Model Top-1 (%)
Vim-Ti (without Position Embedding) 73.0%
Vim-Ti (base) 74.7%
Vim-Ti-F 74.2%
Vim-Ti-F + Position Embedding 75.1%
Vim-Ti-F(H) 76.0%

### v2

@@ IV-D Vim-F(H) Block
PVTv2-B0 [12] 3 0.6 70.5
DeiT-Ti [13] 6 1.3 74.4
MobileViT-XS [14] 2 1.0 74.8
LVT [15] 6 0.9 74.8
Vim-Ti [1] 7 1.5 76.1
LocalVim-T [3] 8 1.5 76.5
Vim-Ti-F (ours) 8 1.8 76.7
PlainMamba-L1 [9] 7 3.0 77.9
RegNetY-1.6G [16] 11 1.6 78.0
Vim-Ti-F(H) (ours) 9 2.1 78.2
MobileViT-S [14] 6 2.0 78.4
EfficientVMamba-S [4] 11 1.3 78.7
DeiT-S [13] 22 4.6 79.8
RegNetY-4G [16] 21 4.0 80.0
Vim-S [1] 26 5.1 80.5
Vim-S-F (ours) 28 6.1 80.9
LocalVim-S[3] 28 4.8 81.0
Swin-T [17] 29 4.5 81.3
Vim-S-F(H) (ours) 37 6.8 81.6
EfficientVMamba-B[4] 33 4.0 81.8
VMamba-T [2] 22 5.6 82.2
V Experiments
This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection and instance segmentation. The specific architectures of Vim-F and Vim-F(H) are as follows: We replaced only the first 6 Vim blocks (25% of the total number of blocks) with either Vim-F or Vim-F(H) blocks to balance performance and efficiency while maintaining approximately the same number of parameters. The number of linear attention heads in Vim-F(H) is set to 24.
V-A Image Classification
Settings. We train the models on ImageNet-1K and evaluate the performance on ImageNet-1K validation set. For fair comparisons, our training settings mainly follow Vim. Due to hardware limitations, our experiments are performed on 3 A6000 GPUs. Therefore, we adjust the total batch size and the initial learning rate to 384 and $3.75\times 10^{-4}$ respectively according to the linear scaling rule[18].
Results. We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table II. Vim-F outperforms the original Vim in terms of performance, and Vim-F(H) further enhances Vim-F’s performance by introducing linear attention. It is worth noting that SOTA ViMs such as VMamba [2], PlainMamba [9], and EfficientVMamba [4] all adopt more complex CNN-Mamba hybrid encoding architectures, and their network structures are similar to advanced convolutional networks like ConvNeXt [19]. Since these works have not conducted the necessary comparative experiments to address this issue, it is reasonable to suspect that the excellent performance of these ViMs may primarily stem from the use of depthwise separable convolutions. However, Vim-F still uses a pure SSM encoder to model visual information. LocalVim[3] is one of the representative works that improve the scanning strategy for Vim. LocalVim has designed 8 scanning strategies and employs neural architecture search to determine the scanning strategy used by each of the 4 scanning branches in every LocalVim block. Additionally, it utilizes a complex spatial and channel attention module (SCAttn) to weight the channels and tokens in each branch feature. Moreover, the total number of blocks in LocalVim has not followed Vim’s 24 but has been reduced to 20. In comparison, our works achieve a better balance between performance, efficiency, and model complexity.
TABLE III: Object detection and instance segmentation results on COCO.
@@ V-A Image Classification
V-B Object Detection and Instance Segmentation
@@
Settings. We use Mask-RCNN as the detector to evaluate the performance of the proposed Vim-F for object detection and instance segmentation on the MSCOCO 2017 dataset. Since the backbone of Vim-F is non-hierarchical, following ViTDet[20], we only used the last feature map from the backbone and generated multi-scale feature maps through a set of convolutions or deconvolutions to adapt to the detector. The remaining settings were consistent with Swin[17]. Specifically, we employ the AdamW optimizer and fine-tune the pre-trained classification models (on ImageNet-1K) for both 12 epochs (1$\times$ schedule). The learning rate is initialized at $1\times 10^{-4}$ and is reduced by a factor of 10×10$\times$10 × at the 9th and 11th epochs.
@@
Results. We summarize the comparison results of Vim-F and Vim-F(H) with other backbones in Table III. It can be seen that our Vim-F consistently outperforms Vim. Similar to the results in classification tasks, Vim-F(H) exhibits better performance than Vim-F and is competitive with advanced ViT architectures.
VI Conclusion
@@
One of the main challenges currently faced by ViMs is the destruction of local two-dimensional dependency relationships when spatial data is flattened into one-dimensional tokens due to the non-causal nature of visual data. Consequently, scanning one-dimensional tokens with specific rules can limit the model’s ability to accurately interpret spatial relationships. We observe that scanning in the frequency domain, leveraging the translation invariance and global nature of Fourier transforms, can reduce the inductive bias of scanning strategies while ensuring that the model always maintains a good global receptive field. We adapt a simple approach of adding the spectrogram with the original feature map, allowing the model to scan simultaneously in both the frequency and spatial domains, thereby establishing a unified representation of visual information across both domains. Furthermore, we employ two trainable parameters to dynamically adjust the information intensity in the frequency and spatial domains, aiming to reduce semantic gaps. We also introduce additional local correlations during patch embedding by using overlapping convolutions. Vim-F(H) enhances the performance of Vim-F by introducing linear attention, which mitigates the sequence order sensitivity of the Mamba encoder. Extensive experiments demonstrate that our work effectively improves the performance of Vim, with Vim-F and Vim-F(H) achieving competitive results.